### v2

@@ 2.4 Quasi-steady viral dynamics in spatial models
In the context of spatial models, the different time scale of the processes alone would not justify neglecting viral diffusion: indeed, we will see in the following that significantly different behaviours may arise. Nevertheless, in some biological scenarios it is reasonable to assume that viral particles do not spread far from infected cells, as explained in the Introduction. When this is the case, the infection is mainly driven by cell-to-cell contact and close-range free virions, hence is is acceptable to ignore viral dynamics. All this leads to the equation
$\begin{cases}\partial_{t}u=D\partial_{xx}^{2}u+pu\Bigl(1-\dfrac{u+i}{K}\Bigr)-\dfrac{\tilde{\beta}}{K}ui\\[8.0pt] \partial_{t}i=D\partial_{xx}^{2}i+\dfrac{\tilde{\beta}}{K}ui-qi\end{cases}$ (U2)
@@
$\begin{cases}\partial_{t}u=\dfrac{D}{K}\partial_{x}[u\partial_{x}(u+i)]+pu\Bigl(1-\dfrac{u+i}{K}\Bigr)-\dfrac{\tilde{\beta}}{K}ui\\[8.0pt] \partial_{t}i=\dfrac{D}{K}\partial_{x}[u\partial_{x}(u+i)]+\dfrac{\tilde{\beta}}{K}ui-qi\end{cases}$ (P2)
in the case of pressure-driven movement (analogous to Eq. (P3)). Both these models have been studied previously [morselli23]. Eq. (U2) fall in the category of classical spatial Lotka–Volterra models for prey and predators, which have been widely studied [dunbar84, li15, morozov06, petrovskii05, petrovskii02]. Eq. (P2) replaces the standard diffusion with nonlinear cross-diffusion: similar models have been introduced in several other works [bubbaheleshaw20, carrillo18, gwiazda19, lorenzi17], but to our knowledge the exact formulation of Eq. (P2) has only been presented in our previous work [morselli23].
@@ 3 Results
$p$ maximal duplication rate of uninfected cells $1.87\times 10^{-2}$ [h-1] [ke00]
$q$ death rate of infected cells $4.17\times 10^{-2}$ [h-1] [ganly00]
$D$ diffusion coefficients (undirected movement) $1.88\times 10^{-4}$ [mm2/h] estimate based on [kim06]
$D$ diffusion coefficients (pressure-driven movement) $1.50\times 10^{-3}$ [mm2/h] estimate based on [kim06]
@@
$K^{2\text{D}}$ tissue carrying capacity in two dimensions $10^{4}$ [cells/mm2] [lodish08]
$q_{v}$ virus clearance rate $1.67\times 10^{-1}$ or $1.00$ [h-1] [mok09], model estimate
$\alpha$ viral burst size $580$ or $3500$ [viruses/cells] [workenhe14, chen01]
@@
$D_{v}$ virus diffusion coefficient $1.00\times 10^{-2}$ [mm2/h] [pooladvand21]
$R_{u}$ initial radius of uninfected cells $2.60$ [$m$] [kim06]
$R_{v}$ initial radius of oncolytic virus $0.50$ [$m$] [pooladvand21]
@@
Table 1: Reference parameter values.
@@ 3.1 Details of numerical simulations
with $V_{0}=2.67\times 10^{4}$ in one spatial dimension and $V_{0}=(2.67\times 10^{3})^{2}$ in two spatial dimensions; these values are obtained by appropriately rescaling the one used in Ref. [pooladvand21], which comes from the number of viral particles used in Ref. [kim06]. The same initial conditions are also employed for the agent-based models.
@@
The viral load released by the death of infected cells $\alpha$ depends highly on the type of virus and ranges from the value 157157\;viruses/cells estimated in Ref. [workenhe14] to the value 35003500\;viruses/cells of Ref. [chen01]; this wide variation is among the motivations that led us to consider different values of $\alpha$. The value of Ref. [workenhe14] however appears too small to show an effective infection in our settings, hence in our simulation we use a minimum value of $\alpha$ equal to 580580\;viruses/cells. The clearance rate of the virus is also a parameter that could significantly vary in different situations: in Ref. [mok09], it has been estimated to $1.67\times 10^{-1}\;$h-1 in vitro, but it is reasonable to assume it to be higher in different settings, due for example to the immune response in vivo or other factors that reduce the viral lifespan. We therefore also consider a higher value of 1.001.00\;h-1. We finally remark that an adequate comparison requires the adaptation of the diffusion coefficient to keep the propagation speed constant in the different models (as listed in Table 1 and further explained in Appendix A.1).
All the simulations have been performed in Matlab 2021b with the techniques described in Ref. [morselli23]. For the individual-based models, we used a temporal step $\tau=0.02\;$h and a spatial step of $\delta=0.1\;$mm both for the one-dimensional and the two-dimensional simulations. Since we only need to keep track of the collective fate of cells in the same lattice point, we used the built-in Matlab functions binornd and mnrnd, which compute random arrays according to binomial and multinomial distributions. In order to allow reproducibility, a random seed has been set at the beginning of each new simulation, ranging from one to five (except in the case of Fig. 10, which shows the results of one hundred simulations). In the figures representing a single simulation, only the one with random seed equal to one is shown: we remark that we performed at least five different simulations also in these cases and the resulting stochastic behaviour arises in all of them in a qualitatively similar way (although the exact quantitative features are clearly different).
Eqs. (U3) and (U2) have been solved with a finite difference scheme explicit in time, using the discretisations $\Delta x=0.1$, $\Delta t=10^{-3}$. We used a forward upwind scheme for the nonlinear diffusion of Eqs. (P3) and (P2), following Ref. [leveque07]; the discretisations are again $\Delta x=0.1$, $\Delta t=10^{-3}$. We remark that the simulations of the electronic supplementary material S4 require a refined discretisation $\Delta x=0.01$, $\Delta t=10^{-4}$ to reduce the diffusion due to the numerical method. In the two-dimensional case we exploited the radial symmetry of the equations.
@@ 3.2 Partial treatment success: travelling waves
Figure 1: (a) Theoretical speed of the travelling wave of viral infection invading an uninfected tumour at carrying capacity, computed following the procedure of Ref. [baabdulla23] (as explained in the Appendix A). The parameters employed are the ones given in Table 1, with three values of $D_{v}$; $q_{v}$ and $\alpha$ are varied so that the value of $\tilde{\beta}$ is constantly equal to $1.02\times 10^{-1}\;$h-1. The dashed cyan line shows the speed of the infection wave obtained for Eq. (U2) with the same parameters. (b) Numerical solutions of Eq. (U3) compared with the position of the theoretical front (shown as a dashed red line). The parameters employed are the ones given in Table 1, with $\alpha=580\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1; we also set $R_{u}=L$ to focus on the infection inside an established tumour.
@@
The speed of travelling waves solutions for PDEs with spatial diffusion can be computed using linearisation techniques. In the case of Eq. (U2), infected cells invade regions in which uninfected cells are at carrying capacity at speed $2\sqrt{D(\beta-q)}$ and slow down as the uninfected cell density decreases [morselli23]; the same result can be also obtained in a rigorous way for a similar system of PDEs [dunbar84]. In the case of Eq. (U3), the minimal speed can be found at the intersection of two appropriate manifolds [baabdulla23]: although it is hard to derive an explicit expression, we can easily rely on numerical computations (see Appendix A). Fig. 1 shows the behaviour of the minimal speed for different parameter values. For the sake of simplicity, we restrict our attention to parameter combinations that maintain the ratio $\alpha/q_{v}$ constant: while this is not a biological constraint, it allows to focus our attention to the specific differences between the modelling approaches in comparable settings. As $q_{v}$ grows, the speed of the infection wave converges to the value $2\sqrt{D(\beta-q)}$, meaning that viral dynamics are irrelevant because the virus is cleared very quickly. Instead, the case $q_{v}=0$ corresponds to $\alpha=0$, which means that viral particles only diffuse without any production or decay; travelling waves are not possible in this hypotetical situation, even though this state has no biological meaning. It is also interesting to observe that for very low values of $q_{v}$ the speed is extremely low: in this case, viral dynamics are mostly dictated by spatial diffusion and the effects of the reaction part of the equation is negligible. In the intermediate cases, the speed is either higher or lower than $2\sqrt{D(\beta-q)}$ depending on whether $D_{v}$ is higher or lower than $D$. In the former case, it is interewsting to observe that there is a specific value of $q_{v}$ that maximises the speed.
@@
Figure 2: Comparison in one spatial dimension between numerical simulations of the discrete model with undirected cell movement (solid lines), the numerical solution of Eq. (U2) (dashed lines) and the numerical solution of Eq. (U3) (dotted black lines) at three different times, with the parameters given in Table 1, $\alpha=580\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1. The viral density of the agent-based model is multiplied by $q_{v}/(\alpha q)$ to allow the comparison with cell numbers; note that the viral density for Eq. (U3) is not shown, as it would superimpose with infected cells. For the agent-based model, the densities of uninfected cells, infected cells and virus are represented respectively in blue, red and purple; the numerical solutions of Eq. (U2) are represented using the same colours. The vertical black lines represent the expected positions of the invasion fronts: the dash-dotted line in panel (a) refers to the infected front, whose speed is represented in Fig. 1, and the dashed line in panels (b) and (c) refers to the uninfected front, travelling at speed $2\sqrt{Dp}$. The horizontal solid black lines show the equilibrium of the ODE given by Eq. (2.7). The horizontal dashed yellow line represents the expected uninfected density at the front for Eq. (U2) given by Eq. (3.2) and the horizontal dash-dotted green line shows the analogue quantity in the case of Eq. (U3) (both are only relevant in panel (c)). The results of the agent-based model are averaged over five simulations. The maximum of the cell density axes in panels (b) and (c) corresponds to the maximum over time of this average (which is larger than the carrying capacity); on the other hand, in panel (a), it has been reduced to allow the comparison between infected cells and virus (and, as a consequence, uninfected cells are not shown).
@@
Let us now describe how the different speed of the infection affects the treatment in comparison to the previous approaches. We start by considering the lower values of $\alpha$ and $q_{v}$ from Table 1, as in this case we expect a more significant difference. Fig. 2, along with the video accompanying it (see electronic supplementary material S2), shows an excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in one spatial dimension. Furthermore, the quasi-steady approximation of Eq. (2.3) holds most of the time, even in the agent-based model (despite some very small stochastic fluctuations), as it is evident from Fig. 2a. Nevertheless, the system’s transient behaviour significantly differs from the solution of Eq. (U2). Fig. 2a shows that the low viral decay allows the viral infection to reach the boundary of the tumour faster than in the case of cell-to-cell infection, in line with the higher speed predicted by theoretical results (see the vertical dash-dotted lines); as a consequence, uninfected cells start to decrease their numbers earlier (Fig. 2b). This difference does not significantly affect the system’s overall behaviour: in all the cases, the travelling wave of uninfected cells invades the surrounding area at the speed $2\sqrt{Dp}$ (vertical dashed lines in Fig. 2b-c) and cell densities at the centre of the tumour converge with damped oscillations to the equilibrium of the corresponding ODE (horizontal solid black lines in Fig. 2), given by Eq. (2.6).
@@
The final peak of the uninfected cells is approximately the one that allows the infection to travel at speed $2\sqrt{Dp}$: in the case of Eq. (2.6), this is given by
@@
Let us first consider the lower values of $\alpha$ and $q_{v}$. The behaviour of numerical simulations is very similar to the one observed in the previous case. Let us recall that, in the case of Eq. (P2), central infections remain localised in the initial position, as described in Ref. [morselli23] (see also Subsection 3.3); this is in sharp contrast with the outcome shown in Fig. 3. Indeed, in the present situation, viral particles can reach the outer region of the tumour without any obstacle and the constraints of cell movement do not stop the propagation of the infection. The similarity of this situation to the case of undirected cell movement suggests that it is more appropriate to compare it to Eq. (U2) than to Eq. (P2). Fig. 3a, along with the video accompanying it (see electronic supplementary material S2), shows that the spread of the infection is extremely similar with respect to the previous case, meaning that the viral diffusion affects the dynamics more significantly than cellular movement; small differences can only be observed by a close comparison of the PDE models, which however are totally unnoticeable in the agent-based model. The only relevant differences between undirected movement and standard diffusion appear to be the maximum value reached by uninfected cell density, which is higher in the former situation, and the profile of the final invasion front, which, in the latter case, is compactly supported and propagates to the surrounding tissues slightly faster. Overall, we can conclude that spatial viral dynamics have a more substantial effect on the evolution of the system than cellular spatial dynamics.
Fig. 3c, along with the video accompanying it (see electronic supplementary material S3), shows that also for high $\alpha$ and $q_{v}$ the outcome is very similar to the case of undirected cell movement, which confirms our previous observation that the infective dynamics are affected mainly by viral spatial spread. This is not trivial

### v3

@@ 2.4 Quasi-steady viral dynamics in spatial models
In the context of spatial models, the different time scale of the processes alone would not justify neglecting viral diffusion: indeed, we will see in the following that significantly different behaviours may arise. Nevertheless, in some biological scenarios it is reasonable to assume that viral particles do not spread far from infected cells, as explained in the Introduction. When this is the case, the infection is mainly driven by cell-to-cell contact and close-range free virions, hence it is acceptable to ignore viral dynamics. All this leads to the equation
$\begin{cases}\partial_{t}u=D_{U}\partial_{xx}^{2}u+pu\Bigl(1-\dfrac{u+i}{K}\Bigr)-\dfrac{\tilde{\beta}}{K}ui\\[8.0pt] \partial_{t}i=D_{U}\partial_{xx}^{2}i+\dfrac{\tilde{\beta}}{K}ui-qi\end{cases}$ (U2)
@@
$\begin{cases}\partial_{t}u=\dfrac{D_{P}}{K}\partial_{x}[u\partial_{x}(u+i)]+pu\Bigl(1-\dfrac{u+i}{K}\Bigr)-\dfrac{\tilde{\beta}}{K}ui\\[8.0pt] \partial_{t}i=\dfrac{D_{P}}{K}\partial_{x}[i\partial_{x}(u+i)]+\dfrac{\tilde{\beta}}{K}ui-qi\end{cases}$ (P2)
in the case of pressure-driven movement (analogous to Eq. (P3)). The quasi-steady approximation of the virus can be mathematically recovered in the limit of fast viral dynamics and slow diffusion. The results of Section 3.3 indeed confirm the validity of the assumption in this parameter regime.
Both these models have been studied previously [53]. Eq. (U2) fall in the category of classical spatial Lotka–Volterra models for prey and predators, which have been widely studied [21, 43, 52, 60, 61]. Eq. (P2) replaces the standard diffusion with nonlinear cross-diffusion: similar models have been introduced in several other works [10, 12, 31, 46], but to our knowledge the exact formulation of Eq. (P2) has only been presented in our previous work [53].
@@ 3 Results
$p$ maximal duplication rate of uninfected cells $1.87\times 10^{-2}$ [h-1] [34]
$q$ death rate of infected cells $4.17\times 10^{-2}$ [h-1] [28]
$D_{U}$ diffusion coefficients (undirected movement) $1.88\times 10^{-4}$ [mm2/h] estimate based on [36]
$D_{P}$ diffusion coefficients (pressure-driven movement) $1.50\times 10^{-3}$ [mm2/h] estimate based on [36]
@@
$K^{2\text{D}}$ tissue carrying capacity in two dimensions $10^{4}$ [cells/mm2] [45]
$q_{v}$ virus clearance rate $1.67\times 10^{-1}$ or $1.00$ [h-1] [51], model estimate
$\alpha$ viral burst size $580$ or $3500$ [viruses/cells] [79, 15]
@@
$D_{v}$ virus diffusion coefficient $1.00\times 10^{-2}$ [mm2/h] [63]
$R_{u}$ initial radius of uninfected cells $2.60$ [$m$] [36]
$R_{v}$ initial radius of oncolytic virus $0.50$ [$m$] [63]
@@
Table 1: Reference parameter values. Note that the value of $\tilde{\beta}$ is given by Eq. (2.4) and depends on ${\beta},\alpha$ and $q_{v}$.
@@ 3.1 Details of numerical simulations
with $V_{0}=2.67\times 10^{4}$ in one spatial dimension and $V_{0}=(2.67\times 10^{3})^{2}$ in two spatial dimensions; these values are obtained by appropriately rescaling the one used in Ref. [63], which comes from the number of viral particles used in Ref. [36]. The same initial conditions are also employed for the agent-based models.
@@
The viral load released by the death of infected cells $\alpha$ depends highly on the type of virus and ranges from the value 157157\;viruses/cells estimated in Ref. [79] to the value 35003500\;viruses/cells of Ref. [15]; this wide variation is among the motivations that led us to consider different values of $\alpha$. The value of Ref. [79] however appears too small to show an effective infection in our settings, hence in our simulation we use a minimum value of $\alpha$ equal to 580580\;viruses/cells. The clearance rate of the virus is also a parameter that could significantly vary in different situations: in Ref. [51], it has been estimated to $1.67\times 10^{-1}\;$h-1 in vitro, but it is reasonable to assume it to be higher in different settings, due for example to the immune response in vivo or other factors that reduce the viral lifespan. We therefore also consider a higher value of 1.001.00\;h-1. We finally remark that an adequate comparison requires the adaptation of the diffusion coefficient to keep the propagation speed constant in the different models (as listed in Table 1 and further explained in Appendix A.1).
All the simulations have been performed in Matlab 2021b with the techniques described in Ref. [53]. For the individual-based models, we used a temporal step $\tau=0.02\;$h and a spatial step of $\delta=0.1\;$mm both for the one-dimensional and the two-dimensional simulations. Since we only need to keep track of the collective fate of cells in the same lattice point, we used the built-in Matlab functions binornd and mnrnd, which compute random arrays according to binomial and multinomial distributions. In order to allow reproducibility, a random seed has been set at the beginning of each new simulation, ranging from one to five (except in the case of Fig. 10, which shows the results of one hundred simulations). In the figures representing a single simulation, only the one with random seed equal to one is shown: we remark that we performed at least five different simulations also in these cases and the resulting stochastic behaviour arises in all of them in a qualitatively similar way (although the exact quantitative features are clearly different).
Eqs. (U3) and (U2) have been solved with a finite difference scheme explicit in time, using the discretisations $\Delta x=0.1$, $\Delta t=10^{-3}$. We used a forward upwind scheme for the nonlinear diffusion of Eqs. (P3) and (P2), following Ref. [42]; the discretisations are again $\Delta x=0.1$, $\Delta t=10^{-3}$. We remark that the simulations of the electronic supplementary material S4 require a refined discretisation $\Delta x=0.01$, $\Delta t=10^{-4}$ to reduce the diffusion due to the numerical method. In the two-dimensional case we exploited the radial symmetry of the equations.
@@ 3.2 Partial treatment success: travelling waves
The speed of travelling waves solutions for PDEs with spatial diffusion can be computed using linearisation techniques. In the case of Eq. (U2), infected cells invade regions in which uninfected cells are at carrying capacity at speed $2\sqrt{D_{U}(\tilde{\beta}-q)}$ and slow down as the uninfected cell density decreases [53]; the same result can be also obtained in a rigorous way for a similar system of PDEs [21]. In the case of Eq. (U3), the minimal speed can be found at the intersection of two appropriate manifolds [4]: although it is hard to derive an explicit expression, we can easily rely on numerical computations (see Appendix A). Fig. 1 shows the behaviour of the minimal speed for different parameter values. For the sake of simplicity, we here restrict our attention to parameter combinations that maintain the ratio $\alpha/q_{v}$ constant: while this is not a biological constraint, it allows to focus our attention to the specific differences between the modelling approaches in comparable settings. Other parameter ranges are explored in Fig. 12.
Figure 1: (a) Theoretical speed of the travelling wave of viral infection invading an uninfected tumour at carrying capacity, computed following the procedure of Ref. [4] (as explained in the Appendix A). The parameters employed are the ones given in Table 1, with three values of $D_{v}$; $q_{v}$ and $\alpha$ are varied so that the value of $\tilde{\beta}$ is constantly equal to $1.02\times 10^{-1}\;$h-1. The dashed cyan line shows the speed of the infection wave obtained for Eq. (U2) with the same parameters. (b) Numerical solutions of Eq. (U3) compared with the position of the theoretical front (shown as a dashed red line). The parameters employed are the ones given in Table 1, with $\alpha=580\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1; we also set $u(0,\cdot)=K$ to focus on the infection inside an established tumour.
Fig. 1 shows that, as $q_{v}$ grows, the speed of the infection wave converges to the value $2\sqrt{D_{U}(\tilde{\beta}-q)}$, meaning that viral dynamics are irrelevant because the virus is cleared very quickly. Instead, the case $q_{v}=0$ corresponds to $\alpha=0$, which means that viral particles only diffuse without any production or decay; travelling waves are not possible in this hypothetical situation, even though this state has no biological meaning. It is also interesting to observe that for very low values of $q_{v}$ the speed is extremely low: in this case, viral dynamics are mostly dictated by spatial diffusion and the effects of the reaction part of the equation is negligible. In the intermediate cases, the speed is either higher or lower than $2\sqrt{D_{U}(\tilde{\beta}-q)}$ depending on whether $D_{v}$ is higher or lower than $D_{U}$ (see also Fig. 12b). It is interesting to observe that when $D_{U}$ is high enough there is a specific value of $q_{v}$ that maximises the speed among the parameter sets that maintain the ratio $\alpha/q_{v}$ constant.
@@
Figure 2: Comparison in one spatial dimension between numerical simulations of the discrete model with undirected cell movement (solid lines), the numerical solution of Eq. (U2) (dashed lines) and the numerical solution of Eq. (U3) (dotted black lines) at three different times, with the parameters given in Table 1, $\alpha=580\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1. The viral density of the agent-based model is multiplied by $q_{v}/(\alpha q)$ to allow the comparison with cell numbers; note that the viral density for Eq. (U3) is not shown, as it would superimpose with infected cells. For the agent-based model, the densities of uninfected cells, infected cells and virus are represented respectively in blue, red and purple; the numerical solutions of Eq. (U2) are represented using the same colours. The vertical black lines represent the expected positions of the invasion fronts: the dash-dotted line in panel (a) refers to the infected front, whose speed is depicted in Fig. 1, and the dashed line in panels (b) and (c) refers to the uninfected front, travelling at speed $2\sqrt{D_{U}p}$. The horizontal solid black lines show the equilibrium of the ODE given by Eq. (2.7). The horizontal dashed yellow line represents the expected uninfected density at the front for Eq. (U2) given by Eq. (3.2) and the horizontal dash-dotted green line shows the analogue quantity in the case of Eq. (U3) (both are only relevant in panel (c)). The results of the agent-based model are averaged over five simulations. The maximum of the cell density axes in panels (b) and (c) corresponds to the maximum over time of this average (which is larger than the carrying capacity); on the other hand, in panel (a), it has been reduced to allow the comparison between infected cells and virus (and, as a consequence, uninfected cells are not shown).
@@
Let us now describe how the different speed of the infection affects the treatment in comparison to the previous approaches. We start by considering the lower values of $\alpha$ and $q_{v}$ from Table 1, as in this case we expect a more significant difference. Fig. 2, along with the video accompanying it (see electronic supplementary material S2), shows an excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in one spatial dimension. Furthermore, the quasi-steady approximation of Eq. (2.3) holds most of the time, even in the agent-based model (despite some very small stochastic fluctuations), as it is evident from Fig. 2a. Nevertheless, the system’s transient behaviour significantly differs from the solution of Eq. (U2). Fig. 2a shows that the low viral decay allows the viral infection to reach the boundary of the tumour faster than in the case of cell-to-cell infection, in line with the higher speed predicted by theoretical results (see the vertical dash-dotted lines); as a consequence, uninfected cells start to decrease their numbers earlier (Fig. 2b). This difference does not significantly affect the system’s overall behaviour: in all the cases, the travelling wave of uninfected cells invades the surrounding area at the speed $2\sqrt{D_{U}p}$ (vertical dashed lines in Fig. 2b-c) and cell densities at the centre of the tumour converge with damped oscillations to the equilibrium of the corresponding ODE (horizontal solid black lines in Fig. 2), given by Eq. (2.6).
@@
The final peak of the uninfected cells is approximately the one that allows the infection to travel at speed $2\sqrt{D_{U}p}$: in the case of Eq. (2.6), this is given by
@@
Let us first consider the lower values of $\alpha$ and $q_{v}$. The behaviour of numerical simulations is very similar to the one observed in the previous case. Let us recall that, in the case of Eq. (P2), central infections remain localised in the initial position, as described in Ref. [53] (see also Subsection 3.3); this is in sharp contrast with the outcome shown in Fig. 3. Indeed, in the present situation, viral particles can reach the outer region of the tumour without any obstacles and the constraints of cell movement do not stop the propagation of the infection. The similarity of this situation to the case of undirected cell movement suggests that it is more appropriate to compare it to Eq. (U2) than to Eq. (P2). Fig. 3a, along with the video accompanying it (see electronic supplementary material S2), shows that the spread of the infection is extremely similar with respect to the previous case, meaning that the viral diffusion affects the dynamics more significantly than cellular movement; small differences can only be observed by a close comparison of the PDE models, which however are totally unnoticeable in the agent-based model. The only relevant differences between undirected movement and standard diffusion appear to be the maximum value reached by uninfected cell density, which is higher in the former situation, and the profile of the final invasion front, which, in the latter case, is compactly supported and propagates to the surrounding tissues slightly faster. Overall, we can conclude that spatial viral dynamics have a more substantial effect on the evolution of the system than cellular spatial dynamics.
Fig. 3c, along with the video accompanying it (see electronic supplementary material S3), shows that also for high $\alpha$ and $q_{v}$ the outcome is very similar to the case of undirected cell movement, which confirms our previous observation that the infective dynamics are affected mainly by viral spatial spread. This is not trivial.